# thanks: This work is based on the author’s M.Sc. dissertation (Ref.[1]), but further correcting and expanding upon it

### v1

@@ V Chiral Anomaly
[…]
$=\frac{1}{2\pi^{2}}\varepsilon_{\alpha\nu\beta\mu}k^{\alpha}_{2}k^{\beta}_{1}$
we can derive the well-known Adler–Bell–Jackiw anomaly [72, *belljackiw] (or Carroll–Field–Jackiw-Torsion [74] if the axial vector derives from torsion):
@@
Here, $\left(j_{5}^{\mu}\right)_{(a)}=\frac{i}{3}A^{\mu}A^{\nu}I_{\mu\nu}$. A similar procedure can be applied to the other surviving diagram; however, in this case, one must impose symmetry between vertices instead of vector-current conservation to obtain
@@
This indicates that the anomaly cancellation conditions of the Standard Model are sufficient to guarantee cancellation in this context as well [75, *anomalycancelation]. This insight presents an intriguing analogy to the pion within the weak field approximation by determining decay rates for each process (e.g., for coherent splitting):
@@
It can be observed that $\Gamma\left(S\to\gamma\gamma\right)/\Gamma\left(S\to SS\right)=9e^{4}/\eta^{4}$. Thus, if $\eta<\sqrt{3}e$, there is a preferential decay of the axial field into a pair of photons. Such dynamics may suggest a mechanism for Primordial Magnetogenesis [19, *magnetobouncing] in cosmological models derived from theories incorporating propagating torsion [78, *torsioninflation]. Furthermore, the Eqs. (73) and (74) yield a term
@@
for the (perturbative) effective action, where we adopt a mixed representation featuring a non-local component, expressed as $\Box^{-1}_{xy}=\int d^{4}p/(2\pi)^{4}e^{-ip\cdot(x-y)}1/p^{2}$, along with a 4D Chern-Simons-like term to underscore the topological nature of the component containing the electromagnetic field, which is a gauge field.
@@
where $f=\mathrm{Tr}e^{-\tau\mathcal{H}}$ and $R=\Delta_{M}\int d^{4}x\left(e^{2}F^{2}_{\mu\nu}+\eta^{2}S^{2}_{\mu\nu}+8\eta^{2}m^{2}S^{2}_{\mu}\right)/6$ is the minimal subtraction term for a MT-type subspace. It follows that $j^{\mu}_{5}$ is given by
@@
$\left.-\frac{1}{12M^{2}\pi^{2}}\left(e^{2}F^{2}_{\alpha\beta}+\eta^{2}S^{2}_{\alpha\beta}+8\eta^{2}m^{2}S^{2}_{\mu}\right)\right]+\frac{8\Delta_{M}}{3}\eta^{2}m^{2}S^{\mu}$
@@
since $\Pi_{\mu}S^{\mu}=0$. Notably, $\delta f/\delta S_{\mu}=-\tau\mathrm{Tr}\left[\eta\left(\left\{\gamma^{5},\Pi^{\mu}\right\}+2\eta S^{\mu}-2m\gamma^{\mu}\gamma^{5}\right)e^{-\tau\mathcal{H}}\right]$, allowing us to show a direct consistency of the effective action obtained within this subspace:
@@
where we employed the Proca action in Eq. (3) to establish $\Delta_{M}\int d^{4}x\delta S^{2}_{\alpha\beta}/\delta S^{\mu}=2\left(M^{2}_{S}S^{\mu}-j_{5}^{\mu}\right)$, with the normalization factor given by $\mathcal{N}=24\pi^{2}/\left(24\pi^{2}-\Delta_{M}\right)$. Its divergence results in
@@
We identify contributions akin to Nieh–Yan’s term [80] ($\sim S^{\mu}\partial_{\mu}S^{2}$ in flat contorted spacetime), along with other contributions to the anomaly that encompass intricate expressions containing derivatives of the invariants. Indeed, aside from $\left(\partial_{\mu}j^{\mu}_{5}\right)_{\text{MT}}$, we need not employ the closed form of the Lagrangian obtained for the quasi-static/constant invariants approximation. Equation (82) indicates what contributions to the anomaly can be anticipated even in broader contexts involving varying fields, affirming that the result remains valid. This observation suggests that achieving stability outside of an MT-type subspace may require further considerations.
But this is not the complete story. To link Eqs. (73) and (74) to Eq. (62), we utilize the classical continuity equation $\left(\partial_{\mu}j^{\mu}_{5}\right)_{\text{MT}}=2imj_{5}$, where the latter term is defined as
$j_{5}=\left<\,\overline{\!{\psi}}\gamma^{5}\psi\right>\sim\frac{im}{16\pi^{2}}\int_{0}^{\infty}d\tau\ e^{-\tau M^{2}}\frac{\mathcal{G}}{\Im\cosh{\left(\tau\overline{X}_{+}\right)}}\ \mathrm{tr}\left(\gamma^{5}e^{-\frac{\tau}{2}V_{\mu\nu}\sigma^{\mu\nu}}\right)$ (83)
In this context, we focus on the term pertinent to the previously derived anomaly contributions within an MT-type subspace, under the condition $S_{\mu\nu}F^{\mu\nu}=\tilde{S}^{\mu\nu}F_{\mu\nu}=0$. Consequently, we find that $\mathrm{tr}\left(\gamma^{5}e^{-\frac{\tau}{2}V_{\mu\nu}\sigma^{\mu\nu}}\right)=-4im\Im\cosh{\left(\tau\overline{X}_{+}\right)}$, leading to
@@
One noteworthy consequence of this anomaly is as follows: Let us define $j^{\mu}=\left<\,\overline{\!{\psi}}\gamma^{\mu}\psi\right>=\left<\,\overline{\!{\psi}}\gamma^{\mu}\left({\rm P}_{R}+{\rm P}_{L}\right)\psi\right>=j^{\mu}_{R}+j^{\mu}_{L}$ and $j^{\mu}_{5}=j^{\mu}_{R}-j^{\mu}_{L}$. Since $A_{\mu}$ is a gauge field, its associated anomaly is inherently tied to topology, specifically, it is proportional to the winding number, which is an integer,
@@ VI Phenomenological Applications
In this section, we will exploit the non-unitarity of the effective action to derive the pair production rates. We will then apply these results within a simplified model of baryogenesis as a proof-of-concept. Furthermore, we will discuss potential extensions and implications of our findings in the realms of condensed matter physics and other related fields.
@@ VI.1 Pair Production Rates
Let us begin by defining the Vacuum Transition Probability as $P_{\text{vac}}=\left|\left<A,S\right|\mathcal{S}_{\text{eff}}\left|A,S\right>\right|^{2}$, where $\mathcal{S}_{\text{eff}}=\exp{\left(-iW\left[A,S\right]\right)}$ is the S-matrix constructed from the full effective action. This probability quantifies the likelihood of remaining within the same background field configuration following scattering processes. According to the framework established by Schwinger [3], we observe that the quantity
@@
provides a measure of the probability of exiting this vacuum state, thereby enabling the creation of any number of particles, therefore moving outside the scope of the effective action. This phenomenon can be modeled as a Poisson process [84], where the pair creation rate per unit volume and time is given by $\rho\left[A,S\right]=-2\Im\mathcal{L}$.
@@
$\left.-2\sqrt{\eta m\left|\widetilde{V}^{\mu\nu}S_{\nu}\right|}\ln{\left[1-\sinh{\left(\frac{2\pi}{E^{(R)}_{V}}\sqrt{\eta m\left|\widetilde{V}^{\mu\nu}S_{\nu}\right|}\right)}e^{-\frac{\pi m^{2}}{E_{V}^{(R)}}}\right]}\right\}$
@@
$\left.-2i\sqrt{\eta m\left|\widetilde{V}^{\mu\nu}S_{\nu}\right|}\ln{\left[1-\sin{\left(\frac{2\pi}{E^{(L)}_{V}}\sqrt{\eta m\left|\widetilde{V}^{\mu\nu}S_{\nu}\right|}\right)}e^{-\frac{\pi m^{2}}{E_{V}^{(L)}}}\right]}\right\}$
It is important to note that the last term in our expression is a pure imaginary number. Consequently, if this term dominates, such as when $4\sqrt{\eta m\left|\widetilde{V}^{\mu\nu}S_{\nu}\right|}=E_{V}^{(L)}$, the vacuum state of left-handed fermions is stabilized, resulting in the absence of particle creation for this chirality. In contrast, for right-handed fermions, we observe an inevitable pair production phenomenon induced by a strong generalized electromagnetic field, irrespective of the value of $\left|\widetilde{V}^{\mu\nu}S_{\nu}\right|$. Remarkably, Eq. (89) also demonstrates that the pair production rate of right-handed fermions can be effectively enhanced by adjusting the parameter $\sqrt{\eta m\left|\widetilde{V}^{\mu\nu}S_{\nu}\right|}$.
When $\left|\widetilde{V}^{\mu\nu}S_{\nu}\right|=0$, the particle creation rates reduce to:
@@
Although both chirality sectors can generate particle pairs, the specific field strengths given by $E^{(R/L)}_{V}=eE_{A}\pm\eta E_{S}$ reveal that one sector will require weaker fields to initiate vacuum decay. As a result, particle creation in this sector begins at an earlier energy, resulting in a net production of fermions with a preferred chirality. However, it is noteworthy that in this case, the chirality-asymmetric instability is not as pronounced as observed in the previous scenario.
@@
$\rho_{2}\left[E,B,S\right]=\frac{EB}{4\pi^{2}}\sum_{k=1}^{\infty}\frac{(-1)^{k}e^{-\frac{\pi kM^{2}}{E}}}{k\sinh{\left(\frac{\pi kB}{eE}\right)}}\Re\cos{\left[\pi k\sqrt{1+2i\frac{B}{E}-\frac{\left(B^{2}+4\eta^{2}m^{2}S^{2}\right)}{E^{2}}}\right]}$ (92)
for the condition $\eta S\leq m$. This expression simplifies to the result obtained by Maroto [26] in the regime where $B=E_{S}=0$ and $2\eta mS\leq E$, yielding
@@
as expected. Notably, in the limits where $B^{2}\ll E^{2}$ and $4\eta^{2}m^{2}S^{2}\ll E^{2}$, we can reformulate Eq. (92) as
@@
This formulation opens avenues for exploring scenarios such as the condition where $E_{S}=B_{A}=0$, $E_{A}\gg B_{S}$, and $E_{A}\gg 2\eta mS$, among other configurations. The corresponding vacuum transition probabilities for these rates, when contrasted with Schwinger’s original rate for QED, can be illustrated in FIG.2. It is noteworthy that when present, the axial vector tends to further stabilize the vacuum. Additionally, we can infer that a dynamical axial vector can create particles on its own. However, this effect is invariably accompanied by the term $S^{2}$, unless the axial vector is massless, indicating that its pair creation rates cannot perfectly replicate those of QED if it possesses mass.
@@ VI.2 Baryogenesis and Cosmology
Let us assume an appropriate coarse-graining process that preserves the Markovian approximation [87], allowing the particle creation rates to remain approximately constant at each step. From Eq. (86), it follows that $n_{5,0}=\int_{t_{i}}^{t_{f}}\rho\left[A,S\right]dt=N_{5}/V_{\text{comoving}}$ represents the total number of particles generated per co-moving volume at zero temperature, where $t_{i}$ and $t_{f}$ denote the initial and final instants of the process. At this time, the universe was dominated by radiation [88, *ParticleDataGroup], leading to
@@
We can then leverage the fact that $n_{5}/s=\text{constant}$ to extrapolate the measured baryon-to-photon ratio $\eta_{B}=n_{5}/n_{\gamma}\approx 6.12\times 10^{-10}$, where $s\approx 7.04n_{\gamma}$ today, but was $s\approx(2\pi^{2}/45)g_{\ast}T_{f}^{3}$ during the radiation-dominated epoch. This allows us to find $n_{5}/s\approx 8.6932\times 10^{-11}\sim 10^{-10}$ in the past, leading to $n_{5}\approx(2\pi^{2}g_{\ast}T_{f}^{3}/45)\left(8.6932\times 10^{-11}\right)$. Furthermore, we assume that energy storage oscillates between the invariants of $S_{\mu}$, sharing it evenly on average as follows: $M^{2}_{S}\left<S^{2}\right>/2=-E^{2}_{S}/2=B^{2}_{S}/2=\mathcal{E}/3$. Thus,
@@
This implies that if the process concludes at temperatures in between $T_{f}\sim 10^{-2}M_{S}$ to $10^{-3}M_{S}$, and the temperature variation spans an order of $10^{-1}$ to $10^{1}$, then values ranging from $\eta\sqrt{\left<S^{2}\right>}=\eta\left<{S}\right>_{\text{rms}}\sim 10^{-14}\,GeV$ to $10^{-12}\,GeV$ would suffice to account for the present baryon asymmetry. Notably, these conditions yield expected values smaller than the total mass of neutrinos, thereby demonstrating consistency with the requirement that $m^{2}-\eta^{2}\left<S^{2}\right>\geq 0$ for well-behaved pair creation rates.
The density of these relics will dilute as $\propto a^{-3}$, where $a(T)\propto\left[T^{3}g_{\ast s}(T)\right]^{-1/3}$ is the scale factor. Thus, to obtain the relic density, we can perform
@@
where we denote the mass in Planck units at the conclusion, i.e., $M_{S}=m_{s}M_{\text{Planck}}$. Ultimately, from $M^{2}_{S}\left<S^{2}\right>=2\mathcal{E}/3$, the relic energy density is approximately
@@
This relic energy density is remarkably low, comfortably situated well below current observational limits on Lorentz-violating interactions from astrophysical sources (e.g., $\gamma$-ray polarimetry) and laboratory tests [91, *Kosteleck_2011]. The corresponding value of the root-mean-square axial field strength $\eta\left<S_{\text{relic}}\right>_{\text{rms}}\lesssim 10^{-61}\,{\rm GeV}$ is also consistent with precision tests of CPT invariance. This parametric suppression arises naturally from the dilution factor $\sim(T_{\text{today}}/T_{f})^{3}$ combined with the stipulation that baryogenesis occurs before the electroweak symmetry breaking. In summary, despite the simplicity of this model, more rigorous constructions of such scenarios could indeed be viable explorations for future research.
It is noteworthy that the relic energy can be finely tuned to approximately $10^{-47}\,GeV^{4}$, while maintaining $\eta\left<S\right>_{\text{rms}}$ below the bounds imposed by experimental evidence. This observation indicates that such axial vector may also serve as a viable candidate for Dark Energy, given that massive vector fields can exhibit negative pressure [93]. In fact, the subset of parameter space where $\Pi_{\mu}S^{\mu}\neq 0$ could present intriguing possibilities for such theoretical proposals [94, *DarkProcaConstraints]. Nonetheless, it is essential to conduct a careful stability analysis to evaluate the implications of this hypothesis rigorously.

### v2

@@ V Chiral Anomaly
we can derive the well-known Adler–Bell–Jackiw anomaly [79, *belljackiw] (or Carroll–Field–Jackiw-Torsion [81] if the axial vector derives from torsion):
@@
A similar procedure can be applied to the other surviving diagram; however, in this case, one must impose symmetry between vertices instead of vector-current conservation to obtain
@@
This indicates that the anomaly cancellation conditions of the Standard Model are sufficient to guarantee cancellation in this context as well [82, *anomalycancelation]. This insight presents an intriguing analogy to the pion within the weak field approximation by determining decay rates for each process (e.g., for coherent splitting):
@@
It can be observed that $\Gamma(S\to\gamma\gamma)/\Gamma(S\to SS)=9e^{4}/\eta^{4}$. Thus, if $\eta<\sqrt{3}e$, there is a preferential decay of the axial field into a pair of photons. Such dynamics may suggest a mechanism for Primordial Magnetogenesis [24, *magnetobouncing] in cosmological models derived from theories incorporating propagating torsion [85, *torsioninflation]. Furthermore, the Eqs. (76) and (77) yield a term
@@
for the (perturbative) effective action, where we adopt a mixed representation featuring a non-local component, expressed as $\Box^{-1}_{xy}=(2\pi)^{-4}\int d^{4}p\,e^{-ip\cdot(x-y)}/p^{2}$, along with a 4D Chern-Simons-like term to underscore the topological nature of the component containing the electromagnetic field, which is a gauge field.
@@
where $f=\mathrm{Tr}e^{\tau\mathcal{H}}$ and $R=(\Delta_{M}/6)\,\int d^{4}x(e^{2}F^{2}_{\mu\nu}+\eta^{2}S^{2}_{\mu\nu}-8\eta^{2}m^{2}S^{2}_{\mu})$ is the minimal subtraction term for a MT-type subspace. It follows that $j^{\mu}_{5}$ is given by
@@
$\left.-\frac{1}{12M^{2}\pi^{2}}\left(e^{2}F^{2}_{\alpha\beta}+\eta^{2}S^{2}_{\alpha\beta}-8\eta^{2}m^{2}S^{2}_{\mu}\right)\right]-\frac{8\Delta_{M}}{3}\eta^{2}m^{2}S^{\mu}$
@@
since $\Pi_{\mu}S^{\mu}=0$. Notably, $\delta f/\delta S_{\mu}=\tau\mathrm{Tr}[\eta(\{\gamma^{5},\Pi^{\mu}\}+2\eta S^{\mu}-2m\gamma^{\mu}\gamma^{5})e^{\tau\mathcal{H}}]$, allowing us to show a direct consistency of the effective action obtained within this subspace:
@@
where we employed the Proca action in Eq. (3) to establish $\Delta_{M}\int d^{4}x\,\delta S^{2}_{\alpha\beta}/\delta S^{\mu}=2(M^{2}_{S}S^{\mu}-j_{5}^{\mu})$, with the normalization factor given by $\mathcal{N}=24\pi^{2}/(24\pi^{2}-\Delta_{M})$. Its divergence results in
@@
We identify contributions akin to Nieh–Yan’s term [87] ($\sim S^{\mu}\partial_{\mu}S^{2}$ in flat contorted spacetime), along with other contributions to the anomaly that encompass intricate expressions containing derivatives of the invariants. Indeed, aside from $(\partial_{\mu}j^{\mu}_{5})_{\text{MT}}$, we need not employ the closed form of the Lagrangian obtained for the quasi-static/constant invariants approximation. Equation (85) indicates what contributions to the anomaly can be anticipated even in broader contexts involving varying fields, affirming that the result remains valid. This observation suggests that achieving stability outside of an MT-type subspace may require further considerations.
But this is not the complete story. To link Eqs. (76) and (77) to Eq. (65), we utilize the classical continuity equation $(\partial_{\mu}j^{\mu}_{5})_{\text{MT}}=2imj_{5}$, where the latter term is defined as
$j_{5}=\left<\,\overline{\!{\psi}}\gamma^{5}\psi\right>\sim\frac{im}{16\pi^{2}}\int_{0}^{\infty}d\tau\ e^{-\tau M^{2}}\frac{\mathcal{G}}{\Im\cosh{\left(\tau\overline{X}_{+}\right)}}\ \mathrm{tr}\left(\gamma^{5}e^{\frac{\tau}{2}V_{\mu\nu}\sigma^{\mu\nu}}\right)$ (86)
In this context, we focus on the term pertinent to the previously derived anomaly contributions within an MT-type subspace, under the condition $S_{\mu\nu}F^{\mu\nu}=\tilde{S}^{\mu\nu}F_{\mu\nu}=0$. Consequently, we find that $\mathrm{tr}(\gamma^{5}e^{\frac{\tau}{2}V_{\mu\nu}\sigma^{\mu\nu}})=4im\Im\cosh{(\tau\overline{X}_{+})}$, leading to
@@
One noteworthy consequence of this anomaly is as follows: Let us define $j^{\mu}=\langle\,\overline{\!{\psi}}\gamma^{\mu}\psi\rangle=\langle\,\overline{\!{\psi}}\gamma^{\mu}({\rm P}_{R}+{\rm P}_{L})\psi\rangle=j^{\mu}_{R}+j^{\mu}_{L}$ and $j^{\mu}_{5}=j^{\mu}_{R}-j^{\mu}_{L}$. Since $A_{\mu}$ is a gauge field, its associated anomaly is inherently tied to topology, specifically, it is proportional to the winding number, which is an integer,
@@ VI Phenomenological Applications
In this section, we will exploit the non-unitarity of the effective action to derive the pair production rates and analyze a simplified model of baryogenesis as a proof-of-concept. Furthermore, we will discuss potential extensions and implications of our findings in the realms of condensed matter physics and other related fields.
@@ VI.1 Pair Production Rates
Let us begin by defining the Vacuum Transition Probability as $P_{\text{vac}}=|\left<A,S\right|\mathcal{S}_{\text{eff}}\left|A,S\right>|^{2}$, where $\mathcal{S}_{\text{eff}}=\exp{(-iW[A,S])}$ is the S-matrix constructed from the full effective action. This probability quantifies the likelihood of remaining within the same background field configuration following scattering processes. According to the framework established by Schwinger [3], we observe that the quantity
@@
provides a measure of the probability of exiting this vacuum state, thereby enabling the creation of any number of particles, therefore moving outside the scope of the effective action. This phenomenon can be modeled as a Poisson process [91], where the pair creation rate per unit volume and time is given by $\rho[A,S]=-2\Im\mathcal{L}$.
@@
$\left.-2\sqrt{\eta m\left|\widetilde{V}^{\mu\nu}S_{\nu}\right|}\ln{\left[1-\sinh{\left(\frac{2\pi}{E^{(L)}_{V}}\sqrt{\eta m\left|\widetilde{V}^{\mu\nu}S_{\nu}\right|}\right)}e^{-\frac{\pi m^{2}}{E_{V}^{(L)}}}\right]}\right\}$
@@
$\left.-2i\sqrt{\eta m\left|\widetilde{V}^{\mu\nu}S_{\nu}\right|}\ln{\left[1-\sin{\left(\frac{2\pi}{E^{(R)}_{V}}\sqrt{\eta m\left|\widetilde{V}^{\mu\nu}S_{\nu}\right|}\right)}e^{-\frac{\pi m^{2}}{E_{V}^{(R)}}}\right]}\right\}$
It is important to note that the last term in our expression is a pure imaginary number. Consequently, if this term dominates, such as when $4\sqrt{\eta m|\widetilde{V}^{\mu\nu}S_{\nu}|}=(2n+1)E_{V}^{(R)}$, for $n$ integer, the vacuum state of right-handed fermions is stabilized, resulting in the absence of particle creation for this chirality. In contrast, for left-handed fermions, we observe an inevitable pair production phenomenon induced by a strong generalized electromagnetic field, irrespective of the value of $|\widetilde{V}^{\mu\nu}S_{\nu}|$. Remarkably, Eq. (92) also demonstrates that the pair production rate of left-handed fermions can be effectively enhanced by adjusting the parameter $\sqrt{\eta m|\widetilde{V}^{\mu\nu}S_{\nu}|}$.
When $|\widetilde{V}^{\mu\nu}S_{\nu}|=0$, the particle creation rates reduce to:
@@
Although both chirality sectors can generate particle pairs, the specific field strengths given by $E^{(R/L)}_{V}=eE_{A}\pm\eta E_{S}$ reveal that one sector might require weaker fields to initiate vacuum decay. As a result, particle creation in this sector can begin at an earlier energy, resulting in a net production of fermions with a preferred chirality. However, it is noteworthy that in this case, the chirality-asymmetric instability is not as pronounced as observed in the previous scenario.
@@
$\rho_{2}\left[E,B,S\right]=\frac{EB}{4\pi^{2}}\sum_{k=1}^{\infty}\frac{(-1)^{k}e^{-\frac{\pi kM^{2}}{E}}}{k\sinh{\left(\frac{\pi kB}{eE}\right)}}\Re\cos{\left[\pi k\sqrt{1-2i\frac{B}{E}+\frac{\left(4\eta^{2}m^{2}S^{2}-B^{2}\right)}{E^{2}}}\right]}$ (95)
for the condition $\eta S\leq m$. This expression simplifies when $S_{\mu}$ is spacelike ($S^{2}<0$, e.g., when it is constant and points at the $z$ direction, as originally obtained by Maroto [31]), in the regime where $B=E_{S}=0$ and $2\eta mS\leq E$, yielding
@@
as expected. Notably, in the limits where $B^{2}\ll E^{2}$ and spacelike axial vector with $4\eta^{2}m^{2}S^{2}\ll E^{2}$, we can reformulate Eq. (95) as
@@
This formulation opens avenues for exploring scenarios such as the condition where $E_{S}=B_{A}=0$, $E_{A}\gg B_{S}$, and $E_{A}\gg 2\eta mS$, among other configurations. The corresponding vacuum transition probabilities for these rates, when contrasted with Schwinger’s original rate for QED, can be illustrated in FIG.2. It is noteworthy that when present, the axial vector tends to further stabilize the vacuum. When the axial vector is time-like, it opposes the effect of the magnetic field, whereas a spacelike pseudo-vector might potentially help destabilize the vacuum under the condition $B^{2}+4\eta^{2}m^{2}S^{2}\geq 1$ and by diminishing the effective mass $M$, which may lead to enhancement in particle creation rates.
Additionally, we can infer that a dynamical axial vector can create particles on its own. However, this effect is invariably accompanied by the term $S^{2}$, unless the axial vector is massless, indicating that its pair creation rates cannot perfectly replicate those of QED if it possesses mass.
@@ VI.2 Baryogenesis and Cosmology
Let us assume an appropriate coarse-graining process that preserves the Markovian approximation [94], allowing the particle creation rates to remain approximately constant at each step. From Eq. (89), it follows that $n_{5,0}=\int_{t_{i}}^{t_{f}}\rho[A,S]dt=N_{5}/V_{\text{comoving}}$ represents the total number of particles generated per co-moving volume at zero temperature, where $t_{i}$ and $t_{f}$ denote the initial and final instants of the process. At this time, the universe was dominated by radiation [95, *ParticleDataGroup], leading to
@@
We can then leverage the fact that $n_{5}/s=\text{constant}$ to extrapolate the measured baryon-to-photon ratio $\eta_{B}=n_{5}/n_{\gamma}\approx 6.12\times 10^{-10}$, where $s\approx 7.04n_{\gamma}$ today, but was $s\approx(2\pi^{2}/45)g_{\ast}T_{f}^{3}$ during the radiation-dominated epoch. This allows us to find $n_{5}/s\approx 8.6932\times 10^{-11}\sim 10^{-10}$ in the past, leading to $n_{5}\approx(2\pi^{2}g_{\ast}T_{f}^{3}/45)(8.6932\times 10^{-11})$. Furthermore, we assume that energy storage oscillates between the invariants of $S_{\mu}$, sharing it evenly on average as follows: $M^{2}_{S}\langle S^{2}\rangle/2=-E^{2}_{S}/2=B^{2}_{S}/2=\mathcal{E}/3$. Thus,
@@
This implies that if the process concludes at temperatures in between $T_{f}\sim 10^{-2}M_{S}$ to $10^{-3}M_{S}$, and the temperature variation spans an order of $10^{-1}$ to $10^{1}$, then values ranging from $\eta\sqrt{\langle S^{2}\rangle}=\eta\langle S\rangle_{\text{rms}}\sim 10^{-14}\,GeV$ to $10^{-12}\,GeV$ would suffice to account for the present baryon asymmetry. Notably, these conditions yield expected values smaller than the total mass of neutrinos, thereby demonstrating consistency with the requirement that $m^{2}-\eta^{2}\langle S^{2}\rangle\geq 0$ for well-behaved pair creation rates.
The density of these relics will dilute as $\propto a^{-3}$, where $a(T)\propto[T^{3}g_{\ast s}(T)]^{-1/3}$ is the scale factor. Thus, to obtain the relic density, we can perform
@@
where we denote the mass in Planck units at the conclusion, i.e., $M_{S}=m_{s}M_{\text{Planck}}$. Ultimately, from $M^{2}_{S}\langle S^{2}\rangle=2\mathcal{E}/3$, the relic energy density is approximately
@@
This relic energy density is remarkably low, comfortably situated well below current observational limits on Lorentz-violating interactions from astrophysical sources (e.g., $\gamma$-ray polarimetry) and laboratory tests [98, *Kosteleck_2011]. The corresponding value of the root-mean-square axial field strength $\eta\langle S_{\text{relic}}\rangle_{\text{rms}}\lesssim 10^{-61}\,{\rm GeV}$ is also consistent with precision tests of CPT invariance. This parametric suppression arises naturally from the dilution factor $\sim(T_{\text{today}}/T_{f})^{3}$ combined with the stipulation that baryogenesis occurs before the electroweak symmetry breaking. In summary, despite the simplicity of this model, more rigorous constructions of such scenarios could indeed be viable explorations for future research.
It is noteworthy that the relic energy can be finely tuned to approximately $10^{-47}\,GeV^{4}$, while maintaining $\eta\langle S\rangle_{\text{rms}}$ below the bounds imposed by experimental evidence. This observation indicates that such axial vector may also serve as a viable candidate for Dark Energy, given that massive vector fields can exhibit negative pressure [100]. In fact, the subset of parameter space where $\Pi_{\mu}S^{\mu}\neq 0$ could present intriguing possibilities for such theoretical proposals [101, *DarkProcaConstraints]. Nonetheless, it is essential to conduct a careful stability analysis to evaluate the implications of this hypothesis rigorously.